Find the slope of the polar curve of the cardiod r=-1+sin 0; 0 = 2
Answer #1

giver r= -lt sino Converting into parametric equations 210) Y(0) = 0(o) Coso = 8(o) sino Gitsino) cose Git sino) sino d(o)=slope at 0 - is slope = m dy da O = 1/2 = yo) xco) 2 =) + sin() sin (1/2) - 2 cos cos ( o + 1 - 1 11 E clo o is undefined. S